y1=12x-5 y2=5x+17, and y1 exceeds y2 by 6
Answer by stanbon(75887)

y1=12x-5 y2=5x+17, and y1 exceeds y2 by 6
-----
Equation:
12x-5 - (5x+17) = 6
-----
12x - 5 - 5x - 17 = 6
7x = 28
x = 4
=============
Cheers,
Stan H.
=============